OpenStudy (anonymous):

Why cant x^2+81 be factored?

OpenStudy (anonymous):

with real numbers, several reasons one is that \[x^2+81\geq 81\] so it is never zero if you could factor it, it would have zeros

OpenStudy (anonymous):

not sure what "reason" means in this context, but you can always factor the DIFFERRENCE of two squares, but not the SUM

OpenStudy (anonymous):

of two square

OpenStudy (anonymous):

what do you mean by difference of two squares but not the sum

OpenStudy (anonymous):

Just simply use the quadratic equation

OpenStudy (kl0723):

example (x^2-4) and (X^2+4) difference and sum

OpenStudy (anonymous):

@satellite73 are you saying you can only factor it if its a minus sign and not adding sign?

OpenStudy (kl0723):

is the example given to you was (x^2-81) it could be factored because is a difference of squares but not (x^2+81) since is a sum :)

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhhhhhhhhh that makes a lot of sense not but what would it be if it was x^2-81?

OpenStudy (anonymous):

(x+9) (x-9)?

OpenStudy (kl0723):

(x+9)(x-9) = x^2-9x+9x-81... 9x-9x =0 and cancel leaving you with x^2-81 :)

OpenStudy (anonymous):

thanks that helps a lot!